OpenStudy (anonymous):

Which of the following points is a solution to the following system of equations? x + y = 1 3x – y = 3 A. (3, 0) B. (5, –1) C. (1, 4) D. (1, 0) @Louis_Tomlinson_Girl

OpenStudy (anonymous):

i just going to say this now i'm not very good at this type of math here

OpenStudy (anonymous):

Plug them in, you are given the x and y value and solve.

OpenStudy (anonymous):

i need help @AEYES13

OpenStudy (anonymous):

x + y = 1 3x – y = 3 _____________-- 4x=4 x=1 1+y=1 y=0

OpenStudy (anonymous):

The answer is D. You don't even have to know how to do systems of equations, just plug in (x,y) into each equation until you get the correct pair of solutions.

OpenStudy (anonymous):

Yep !